# Discovering Artificial Viscosity Models for Discontinuous Galerkin approximation of Conservation Laws using Physics-Informed Machine Learning

### v1

@@ Abstract
  • Finite element-based high-order solvers of conservation laws offer large accuracy but face challenges near discontinuities due to the Gibbs phenomenon. Artificial viscosity is a popular and effective solution to this problem based on physical insight. In this work, we present a physics-informed machine learning algorithm to automate the discovery of artificial viscosity models in a non-supervised paradigm. The algorithm is inspired by reinforcement learning and trains a neural network acting cell-by-cell (the viscosity model) by minimizing a loss defined as the difference with respect to a reference solution thanks to automatic differentiation. This enables a dataset-free training procedure. We prove that the algorithm is effective by integrating it into a state-of-the-art Runge-Kutta discontinuous Galerkin solver. We showcase several numerical tests on scalar and vectorial problems, such as Burgers’ and Euler’s equations in one and two dimensions. Results demonstrate that the proposed approach trains a model that is able to outperform classical viscosity models. Moreover, we show that the learnt artificial viscosity model is able to generalize across different problems and parameters.
@@ 1 Introduction
Our goal in this paper is to propose and test an algorithm to train a NN-based artificial viscosity model in a non-supervised learning framework. To this end, we formulate the problem as an RL task and we enhance it with physics-informed machine learning (PIML) [32, 26]. Namely, thanks to automatic differentiation, we can differentiate through the RL environment, effectively embedding information on the physics of the problem into the learning dynamics. The resulting approach shares some similarities with model-based RL such as [25], differentiable RL [42], and optimal control. Therefore, our approach overcomes the shortcomings of supervised learning and is cheaper than RL since we can exploit the differentiability of the environment (the DG solver). The reward is defined as the dissimilarity from a reference solution with the addition of suitable regularization terms. This also enables seamless integration of noisy data obtained from measurements of quantities of interest into the loss function, thereby being able to learn new physical models directly from data. The NN that we employ is a generalization of the one presented in [16] aimed at increasing flexibility and accuracy. An ablation study proves the efficacy of the proposed modifications.
@@ 2.1 The semi-discrete discontinuous Galerkin formulation
For the spatial discretization, we employ the discontinuous Galerkin method. Let $\mathcal{K}_{h}$ be a shape-regular family of open, disjoint cells (intervals for $d=1$ or triangles for $d=2$) $K$ that approximates $\Omega$. Moreover, let $\mathcal{E}_{h}$ be the set of all faces (points for $d=1$ or edges for $d=2$). On each face $F\in\mathcal{E}_{h}$, approximating the boundary $\partial\Omega$, a single type of boundary condition is imposed at a time (e.g. only Dirichlet or Neumann). We define $h=\max_{K\in\mathcal{K}_{h}}h_{K}$, where $h_{K}$ denotes the diameter of $K$. We define the averages on and jumps across the edges for a smooth enough scalar function $v:\Omega\to\mathbb{R}$ and for a smooth enough vector-valued function $\boldsymbol{q}:\Omega\to\mathbb{R}^{n}$. Let us consider two neighboring elements $K^{+}$ and $K^{-}$, let $\boldsymbol{n}^{+}$ and $\boldsymbol{n}^{-}$ be their outward normal unit vector and let $v^{\pm},\boldsymbol{q}^{\pm}$ be the restrictions to $K^{\pm}$, respectively. In accordance with [3] we define
@@
    $\{\!\!\{v\}\!\!\}=\frac{1}{2}(v^{-}+v^{+}),\qquad\{\!\!\{\boldsymbol{q}\}\!\!\}=\frac{1}{2}(\boldsymbol{q}^{-}+\boldsymbol{q}^{+}),\qquad[\![v]\!]=v^{+}\boldsymbol{n}^{+}+v^{-}\boldsymbol{n}^{-},\qquad[\![\boldsymbol{q}]\!]=\boldsymbol{q}^{+}\cdot\boldsymbol{n}^{+}+\boldsymbol{q}^{-}\cdot\boldsymbol{n}^{-}.$
  • on boundary edges, defined the symmetric outer product $\boldsymbol{q}\odot\boldsymbol{n}=(\boldsymbol{v}\otimes\boldsymbol{n}+\boldsymbol{n}\otimes\boldsymbol{v})$
    $\{\!\!\{v\}\!\!\}=v,\qquad\{\!\!\{\boldsymbol{q}\}\!\!\}=\boldsymbol{q},\qquad[\![v]\!]=v\boldsymbol{n},\qquad[\![\boldsymbol{q}]\!]=\boldsymbol{q}\odot\boldsymbol{n}.$
To further stabilize the scheme, as shown in [8, 31], we recall the following weighted average, which introduces a stabilizing numerical flux
$\{\!\!\{\boldsymbol{q}\}\!\!\}_{\alpha}=\{\!\!\{\boldsymbol{f}(\boldsymbol{q})\}\!\!\}+\frac{[\![\alpha]\!]}{2}[\![\boldsymbol{q}]\!]\quad\forall\;F\in\mathcal{E}_{h},$
where $\alpha$ represents a pair of scalars $\alpha^{\pm}$ associated with the elements $K^{\pm}$. For some choices of $\alpha$ we can recover already well-known numerical fluxes: if $\boldsymbol{f}$ is the identity and $\alpha^{\pm}=\frac{1}{2}\pm\frac{1}{2}$ we have the upwind flux [4], if $\alpha$ is the largest eigenvalue (in absolute value) $\lambda$ of the Jacobian $\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}$ we have the Rusanov flux [52]. To construct the semi-discrete formulation, we define for any integer $k\geq 1$ the finite element space of discontinuous piecewise polynomial functions as
@@
    $\mathscr{A}_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})$ $=\sum_{K\in\mathcal{K}_{h}}\int_{\Omega}\mu(\boldsymbol{u}_{h})\nabla\boldsymbol{u}_{h}\cdot\nabla\boldsymbol{v}_{h}$
    $+\sum_{F\notin\Gamma_{N}}\frac{\tau k^{2}}{\left|F\right|}\int_{F}\mu(\boldsymbol{u}_{h})[\![\boldsymbol{u}_{h}]\!][\![\boldsymbol{v}_{h}]\!]-\sum_{F\in\mathring{\mathcal{E}}_{h}}\{\!\!\{\nabla\boldsymbol{u}_{h}\}\!\!\}[\![\boldsymbol{v}_{h}]\!]+[\![\boldsymbol{u}_{h}]\!]\{\!\!\{\nabla\boldsymbol{v}_{h}\}\!\!\}\quad\forall\;\boldsymbol{u}_{h},\boldsymbol{v}_{h}\in\boldsymbol{V}_{k}^{h},$
@@
    $\mathscr{B}_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})=\sum_{K\in\mathcal{K}_{h}}\int_{K}\boldsymbol{f}(\boldsymbol{u}_{h})\cdot\nabla\boldsymbol{v}_{h}-\sum_{F\in\mathring{\mathcal{E}}_{h}}\int_{F}\{\!\!\{\boldsymbol{f}(\boldsymbol{u}_{h})\}\!\!\}_{\lambda}\cdot[\![\boldsymbol{v}_{h}]\!]\quad\forall\;\boldsymbol{u}_{h},\boldsymbol{v}_{h}\in\boldsymbol{V}_{k}^{h}.$
@@ 2.3 Fully-discrete formulation
Specifically, we utilize one of the two following options: a third-order Strong Stability Preserving Runge-Kutta (SSPR3) scheme [19] or a fourth-order low-storage explicit Runge-Kutta with five stages [12]. The choice between the two is made so to minimize the number of total stages while maintaining a stable solution.
@@ 2.4 Convective flux models
  • the Euler system, namely, given the ideal gas law for pressure $p=(\gamma-1)(e-\frac{1}{2}\rho\left|v\right|^{2})$, where $\gamma=7/5$ is the heat capacity ratio, $e$ is the energy, $\rho$ is the density and $v$ is the velocity, we have ($m=3$):
@@
    where $\boldsymbol{v}=\begin{bmatrix}v_{1}\\ v_{2}\end{bmatrix}$ is the velocity field, and the pressure $p=(\gamma-1)(e-\frac{1}{2}\rho\left\|\boldsymbol{v}\right\|_{2}^{2})$ is defined analogously as before.
@@ 3.1 An outline of artificial viscosity
    where $c_{\max{}}$ is a scalar parameter to be tuned for each problem [48, 34] and $\boldsymbol{f}^{\prime}$ defines the velocity of propagation of the problem (for $m>1$ we consider the restriction to the equation of the representative variable). Notice that the definition of $\mu_{\max{}}$ assures that the viscous contribution is at most linear.
@@ 3.3 Basic concepts of Neural networks
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{i=1}^{N_{\textnormal{train}}}\left\|\hat{\boldsymbol{y}}-\boldsymbol{y}(\boldsymbol{x};\boldsymbol{\theta})\right\|_{p},$ (13)
where $\left\|\cdot\right\|_{p}$ indicates the discrete $p$-norm, $p=1,2$. A sequence of gradient updates made by iterating over the dataset one time is called an epoch. Even if we will propose an alternative paradigm, this is what is used in literature to train the artificial viscosity model and it is useful to fix this as a benchmark to be compared to our approach. Regardless of the paradigm used to train the NN, if properly optimized, neural networks are able to generalize, meaning that they make good predictions even for input data not used during the training phase. For a comprehensive description of neural networks, we direct the reader to [22, 63].

### v2

@@ Abstract
  • Finite element-based high-order solvers of conservation laws offer large accuracy but face challenges near discontinuities due to the Gibbs phenomenon. Artificial viscosity is a popular and effective solution to this problem based on physical insight. In this work, we present a physics-informed machine learning algorithm to automate the discovery of artificial viscosity models. We refer to the proposed approach as an “hybrid” approach which stands at the edge between supervised and unsupervised learning. More precisely, the proposed “hybrid” paradigm is not supervised in the classical sense as it does not utilize labeled data in the traditional way but relies on the intrinsic properties of the reference solution. The algorithm is inspired by reinforcement learning and trains a neural network acting cell-by-cell (the viscosity model) by minimizing a loss defined as the difference with respect to a reference solution thanks to automatic differentiation. This enables a dataset-free training procedure. We prove that the algorithm is effective by integrating it into a state-of-the-art Runge-Kutta discontinuous Galerkin solver. We showcase several numerical tests on scalar and vectorial problems, such as Burgers’ and Euler’s equations in one and two dimensions. Results demonstrate that the proposed approach trains a model that is able to outperform classical viscosity models. Moreover, we show that the learnt artificial viscosity model is able to generalize across different problems and parameters.
@@ 1 Introduction
Our goal in this paper is to propose and test an algorithm to train a NN-based artificial viscosity model in a hybrid learning framework, which stands at the edge between supervised and unsupervised learning. To this end, we formulate the problem as an RL task and we enhance it with physics-informed machine learning (PIML) [32, 26]. Namely, thanks to automatic differentiation, we can differentiate through the RL environment, effectively embedding information on the physics of the problem into the learning dynamics. The resulting approach shares some similarities with model-based RL such as [25], differentiable RL [42], and optimal control. Therefore, our approach overcomes the shortcomings of supervised learning and is cheaper than RL since we can exploit the differentiability of the environment (the DG solver). The reward is defined as the dissimilarity from a reference solution with the addition of suitable regularization terms. This also enables seamless integration of noisy data obtained from measurements of quantities of interest into the loss function, thereby being able to learn new physical models directly from data. The NN that we employ is a generalization of the one presented in [16] aimed at increasing flexibility and accuracy. An ablation study proves the efficacy of the proposed modifications.
@@ 2.1 The semi-discrete discontinuous Galerkin formulation
For the spatial discretization, we employ the discontinuous Galerkin method. Let $\mathcal{K}_{h}$ be a shape-regular family of open, disjoint cells (intervals for $d=1$ or triangles for $d=2$) $K$ that approximates $\Omega$. Moreover, let $\mathcal{E}_{h}$ be the set of all faces (points for $d=1$ or edges for $d=2$). On each face $F\in\mathcal{E}_{h}$, approximating the boundary $\partial\Omega$, a single type of boundary condition is imposed at a time (e.g. only Dirichlet or Neumann). We define $h=\max_{K\in\mathcal{K}_{h}}h_{K}$, where $h_{K}$ denotes the diameter of $K$. We define the averages on and the jumps across the edges for a smooth enough scalar function $v:\Omega\to\mathbb{R}$, for a smooth enough vector-valued function $\boldsymbol{q}:\Omega\to\mathbb{R}^{n}$ and a smooth enough tensor-valued function $\boldsymbol{\psi}$. Let us consider two neighboring elements $K^{+}$ and $K^{-}$, let $\boldsymbol{n}^{+}$ and $\boldsymbol{n}^{-}$ be their outward normal unit vector and let $v^{\pm},\boldsymbol{q}^{\pm}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\boldsymbol{\psi}^{\pm}}$ be the restrictions to $K^{\pm}$, respectively. In accordance with [3], defined the symmetric outer product $\boldsymbol{q}\odot\boldsymbol{n}=\frac{1}{2}(\boldsymbol{v}\otimes\boldsymbol{n}+\boldsymbol{n}\otimes\boldsymbol{v})$, we define
@@
    $\{\!\!\{v\}\!\!\}=\frac{1}{2}(v^{-}+v^{+}),\qquad[\![v]\!]=v^{+}\boldsymbol{n}^{+}+v^{-}\boldsymbol{n}^{-},$
    $\{\!\!\{\boldsymbol{q}\}\!\!\}=\frac{1}{2}(\boldsymbol{q}^{-}+\boldsymbol{q}^{+}),\qquad[\![\boldsymbol{q}]\!]=\boldsymbol{q}^{+}\odot\boldsymbol{n}^{+}+\boldsymbol{q}^{-}\odot\boldsymbol{n}^{-},$
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\!\!\{\boldsymbol{\psi}\}\!\!\}=\frac{1}{2}(\boldsymbol{\psi}^{-}+\boldsymbol{\psi}^{+}),}\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[\![\boldsymbol{\psi}]\!]=\boldsymbol{\psi}^{-}\boldsymbol{n}^{-}+\boldsymbol{\psi}^{+}\boldsymbol{n}^{+}.}$
  • on boundary edges:
    $\{\!\!\{v\}\!\!\}=v,\quad\{\!\!\{\boldsymbol{q}\}\!\!\}=\boldsymbol{q},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\!\!\{\boldsymbol{\psi}\}\!\!\}=\boldsymbol{\psi},\quad}$
    $[\![v]\!]=v\boldsymbol{n},\quad[\![\boldsymbol{q}]\!]=\boldsymbol{q}\odot\boldsymbol{n},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[\![\boldsymbol{\psi}]\!]=\boldsymbol{\psi}\boldsymbol{n}}.$
To further stabilize the scheme, as shown in [8, 31], we introduce the following weighted average, which introduces a stabilizing numerical flux
$\{\!\!\{\boldsymbol{q}\}\!\!\}_{\alpha}=\{\!\!\{\boldsymbol{f}(\boldsymbol{q})\}\!\!\}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha}}{2}[\![\boldsymbol{q}]\!]\quad\forall\;F\in\mathcal{E}_{h},$ (2)
where $\alpha\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}>0}$ is a scalar. For some choices of $\alpha$ we can recover already well-known numerical fluxes: if $\boldsymbol{f}$ is the identity and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha=1}$ we have the upwind flux [4], if $\alpha$ is the largest eigenvalue (in absolute value) $\lambda$ of the Jacobian $\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}$ we have the Rusanov flux [52]. To construct the semi-discrete formulation, we define for any integer $k\geq 1$ the finite element space of discontinuous piecewise polynomial functions as
@@
    $\mathscr{A}_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})$ $=\sum_{K\in\mathcal{K}_{h}}\int_{\Omega}\mu(\boldsymbol{u}_{h})\nabla\boldsymbol{u}_{h}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}\nabla\boldsymbol{v}_{h}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}\Omega}$
    $+\sum_{F\notin\Gamma_{N}}\frac{\tau k^{2}}{\left|F\right|}\int_{F}\mu(\boldsymbol{u}_{h})[\![\boldsymbol{u}_{h}]\!]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}[\![\boldsymbol{v}_{h}]\!]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}\Sigma}$
    $-\sum_{F\in\mathring{\mathcal{E}}_{h}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{F}}\{\!\!\{\nabla\boldsymbol{u}_{h}\}\!\!\}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}[\![\boldsymbol{v}_{h}]\!]+[\![\boldsymbol{u}_{h}]\!]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}\{\!\!\{\nabla\boldsymbol{v}_{h}\}\!\!\}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}\Sigma}\quad\forall\;\boldsymbol{u}_{h},\boldsymbol{v}_{h}\in\boldsymbol{V}_{k}^{h},$
@@
    $\mathscr{B}_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})$ $=\sum_{K\in\mathcal{K}_{h}}\int_{K}\boldsymbol{f}(\boldsymbol{u}_{h}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}\nabla\boldsymbol{v}_{h}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}\Omega}$
    $-\sum_{F\in\mathring{\mathcal{E}}_{h}}\int_{F}\{\!\!\{\boldsymbol{u}_{h}\}\!\!\}_{\lambda}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:}[\![\boldsymbol{v}_{h}]\!]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}\Sigma}\quad\forall\;\boldsymbol{u}_{h},\boldsymbol{v}_{h}\in\boldsymbol{V}_{k}^{h},$
    where $\{\!\!\{\boldsymbol{u}_{h}\}\!\!\}_{\lambda}$ indicates the Rusanov flux, as defined in Eq. (2).
@@ 2.3 Fully-discrete formulation
Specifically, we utilize one of the two following options: a third-order Strong Stability Preserving Runge-Kutta (SSPR3) scheme [19] or a fourth-order low-storage explicit Runge-Kutta with five stages [12]. The choice between the two is made so to minimize the number of total stages while maintaining a stable solution. To further verify the consistency of the results, we compute the maximum Courant number
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C=\max_{\boldsymbol{x}\in\Omega}\left|\boldsymbol{f}^{\prime}(\boldsymbol{u}_{h}(\boldsymbol{x},t))\right|\frac{\Delta t}{h}.}$
@@ 2.4 Convective flux models
  • the Euler system, namely, given the ideal gas law for pressure $p=(\gamma-1)(e-\frac{1}{2}\rho\left|v\right|^{2})$, where $\gamma=7/5$ is the heat capacity ratio, $e$ is the total energy, $\rho$ is the density and $v$ is the velocity, we have ($m=3$):
@@
    where $\boldsymbol{v}=\begin{bmatrix}v_{1}\\ v_{2}\end{bmatrix}$ is the velocity field, $e$ is the total energy, and the pressure $p=(\gamma-1)(e-\frac{1}{2}\rho\left\|\boldsymbol{v}\right\|_{2}^{2})$ is defined analogously as before.
@@ 3.1 An outline of artificial viscosity
    where $c_{\max{}}$ is a scalar parameter to be tuned for each problem [48, 34] and $\boldsymbol{f}^{\prime}$ defines the velocity of propagation of the problem. For $m>1$ we consider the restriction to the equation of the representative variable. This restriction does not constitute a limitation; all the components of $\boldsymbol{f}^{\prime}$ can be utilized as inputs for the neural network. The primary drawback of this approach is the increased computational cost. Notice that the definition of $\mu_{\max{}}$ assures that the viscous contribution is at most linear.
@@ 3.3 Basic concepts of Neural networks
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{i=1}^{N_{\textnormal{train}}}\left\|\hat{\boldsymbol{y}}-\boldsymbol{y}(\boldsymbol{x};\boldsymbol{\theta})\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}},$ (14)
where $\left\|\cdot\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ indicates the discrete ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}$-norm, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}=1,2$. A sequence of gradient updates made by iterating over the dataset one time is called an epoch. Even if we will propose an alternative paradigm, this is what is used in literature to train the artificial viscosity model and it is useful to fix this as a benchmark to be compared to our approach. Regardless of the paradigm used to train the NN, if properly optimized, neural networks are able to generalize, meaning that they make good predictions even for input data not used during the training phase. For a comprehensive description of neural networks, we direct the reader to [22, 63].